# Mamba-based Efficient Spatio-Frequency Motion Perception for Video Camouflaged Object Detection

### v1

@@ Abstract
Existing video camouflaged object detection (VCOD) methods primarily rely on spatial appearance features to perceive motion cues for breaking camouflage. However, the high similarity between foreground and background in VCOD results in limited discriminability of spatial appearance features (e.g., color and texture), restricting detection accuracy and completeness. Recent studies demonstrate that frequency features can not only enhance feature representation to compensate for appearance limitations but also perceive motion through dynamic variations in frequency energy. Furthermore, the emerging state space model called Mamba, enables efficient perception of motion cues in frame sequences due to its linear-time long-sequence modeling capability. Motivated by this, we propose a novel visual camouflage Mamba (Vcamba) based on spatio-frequency motion perception that integrates frequency and spatial features for efficient and accurate VCOD. Specifically, we propose a receptive field visual state space (RFVSS) module to extract multi-scale spatial features after sequence modeling. For frequency learning, we introduce an adaptive frequency component enhancement (AFE) module with a novel frequency-domain sequential scanning strategy to maintain semantic consistency. Then we propose a space-based long-range motion perception (SLMP) module and a frequency-based long-range motion perception (FLMP) module to model spatio-temporal and frequency-temporal sequences in spatial and frequency phase domains. Finally, the space and frequency motion fusion module (SFMF) integrates dual-domain features for unified motion representation. Experimental results show that our Vcamba outperforms state-of-the-art methods across 6 evaluation metrics on 2 datasets with lower computation cost, confirming the superiority of Vcamba. Our code is available at: https://github.com/BoydeLi/Vcamba.
Index Terms:
Video camouflaged object detection, frequency learning, motion perception, visual Mamba.
@@ I Introduction
However, it is challenging for both humans and machines to accurately localize camouflaged objects based solely on appearance in static scenes. In nature, organisms identify prey or predators by perceiving dynamic changes in the environment, as motion reveals inconsistencies with the static background. This phenomenon aligns with the two-stream hypothesis of neuroscience [18], which indicates that motion is crucial for breaking camouflage. Based on this principle, existing video camouflaged object detection (VCOD) methods [19, 20, 21, 22] break camouflage by explicitly modeling motion cues in video sequences like optical flow or implicitly learning them through temporal context.
@@
Although existing methods have made progress, relying solely on spatial appearance features limits performance. The high similarity between foreground and background in texture, color, and pattern results in limited discriminative cues and increased background noise in spatial domain. Recent COD methods [23, 24, 25] have demonstrated the effectiveness of frequency features for enhancing foreground-background distinction. More critically, spatial features are limited by local pixel dependencies and lack global semantic awareness, whereas frequency features inherently provide global representation [26, 27]. Moreover, previous studies[28, 29] have shown that inter-frame motion can be effectively represented by the dynamic variations in frequency-domain energy. Therefore, Integrating frequency features into VCOD supplements discriminative cues and enables global motion perception, overcoming spatial feature constraints. As shown in Fig. 1, frequency features enables our model to deliver finer detail segmentation and more accurate global localization.
@@
In addition, most existing COD and VCOD methods [13, 21, 23] are based on convolutional neural networks (CNNs) and Transformers. However, CNN-based methods struggle to capture global features due to local receptive fields, while Transformer-based methods enable global modeling through self-attention, but facing efficiency issues from quadratic complexity. Since VCOD requires both global perception and high real-time performance, the emerging state space model called Mamba [30] enables global modeling in linear-time, making it an ideal fit for VCOD. It also further improves computational efficiency via hardware-optimized techniques such as parallel scanning. Mamba has been extended from natural language processing (NLP) to vision tasks for efficient long-range modeling of images after converting them into 1D representations [31, 32]. It achieves strong performance in downstream visual tasks [32, 33, 34].
Based on selective scan, Mamba can dynamically adjust state update parameters to adaptively focus on critical regions based on input [29]. So it enables adaptive distinction of foreground–background patches and frequency component learning. Meanwhile, inter-frame motion perception inherently involves long-range dependency modeling of frame sequences. Based on all these insights, we propose Vcamba, a Mamba-based spatio-frequency motion perception network.
Specifically, inspired by the traditional receptive field block (RFB), we propose a receptive field visual state space (RFVSS) block by embedding multi-scale receptive fields into the VssBlock of VMamba [32] to complement the global dependencies with fine-grained local details. In addition, we design a novel frequency-domain sequential scanning strategy that unfolds frequency components in both low-to-high and high-to-low orders, addressing the semantic incoherence caused by large intensity disparities between frequency components. Then we perform long-range dependency modeling across frequency component sequence to implement adaptive frequency enhancement (AFE) for COD. We propose a dual-branch framework to perceive motion from spatio-temporal and frequency-temporal sequences, comprising a space-based long-range motion perception (SLMP) and a frequency-based long-range motion perception (FLMP) module. By constructing spatio-temporal and frequency-temporal maps, SLMP and FLMP adopt novel spatio-temporal and frequency-temporal scanning strategies to model long-range temporal dependencies, enabling implicit motion perception in the spatial and frequency domains. We further explore the role of frequency-domain properties in COD. As shown in Fig. 1, we apply fast fourier transform (FFT) to RGB images to obtain amplitude and phase maps, which are then reconstructed into spatial domain via inverse FFT (iFFT). Visualization results show that the amplitude map reflects energy distribution and texture details, while the phase map encodes image layout and object structures. In FLMP, we only use the phase map as motion perception carrier, modeling the temporal variations of object structures, which effectively suppresses texture and color distractions in COD. Furthermore, to alleviate domain and structural discrepancies between spatial and frequency motion features, we propose a sequence-to-sequence and point-to-point dual-domain sequence concatenation strategy for global and local spatial-frequency motion fusion (SFMF). Finally, the unified motion representations are decoded to enable accurate segmentation. The main contributions are as follows:
@@
  • We propose a frequency-domain sequential scanning strategy that unfolds FFT-derived frequency components in both low-to-high and high-to-low spiral order to achieve adaptive frequency component enhancement.
@@
  • Our Vcamba achieves state-of-the-art performance across six evaluation metrics on two benchmark datasets, and has lower computation cost compared to other VCOD methods, demonstrating the effectiveness and high efficiency of our method.
@@ II-A Video Camouflaged Object Detection
Existing COD methods [35, 36, 25, 17] use CNNs and transformers to extract static visual features like color and texture to break camouflage. Fan et al. [35] first introduced COD task and established a two-stage search-and-identify network as a baseline. However, static visual features are easily disrupted by camouflage, making them less discriminative. In contrast, dynamic motion cues across video frames are considered crucial for breaking camouflage. Traditional video methods exploit motion cues [37] or optical flow [38] to perceive camouflaged objects. Cheng et al. [19] argued that the ambiguous and indistinct boundaries of camouflaged objects hinder the accuracy of explicit motion estimation methods such as optical flow. To address this, they proposed a Transformer-based network called SLTNet that implicitly captures motion through long-range modeling. Hui et al. [21] integrated implicit and explicit motion learning into a unified framework, proposing IMEX, which leverages global frame alignment and explicit motion supervision. Then, they further constructed self-prompts to harness the powerful Segment Anything Model for precise segmentation in [20]. Most of the above methods explore motion cues in the spatial domain. However, discriminative features in the spatial domain are limited due to the similarity between foreground and background. We further explore frequency feature and integrate it with spatial motion cues to achieve complementary enhancement.
@@ II-B Visual Mamba
Mamba [39] is a general sequence modeling architecture based on the selective State Space Model (SSM). Its SSM effectively captures long-range dependencies, demonstrating linear time complexity relative to sequence length. This addresses the limitations of Transformers in long-range dependency modeling performance and computational efficiency. Mamba was initially proposed in NLP for language sequences modeling [30]. Then, the first visual Mamba backbone, Vision Mamba [31], was introduced. It utilizes a bidirectional scanning method to serialize images into 1D sequences, with the SSM capturing long-range dependencies in the image sequence. Based on this, Liu et al. [32] proposed the more powerful Vmamba, which performs four-directional scanning in both horizontal and vertical directions. Furthermore, to better adapt Visual Mamba to downstream salient object detection, He et al. [40] proposed a saliency Mamba called Samba, which employs a spatial neighboring scanning strategy to preserve the spatial continuity of salient objects. In addition, Mamba has also been widely applied to image classification [33], semantic segmentation [32], and multi-modal learning [34]. However, the feasibility of Mamba in COD remains under-explored. Considering that Mamba can adaptively focus on critical regions by its dynamical state parameter update, we utilize it to mine camouflaged object regions and adaptively enhance relevant frequency components. Meanwhile, we use its long-range dependency modeling to effectively and efficiently perceive motion cues across frame sequences.
@@ II-C Frequency Learning
Frequency learning is widely used in signal analysis to reveal the time-frequency characteristics and energy distribution patterns of signals. The rich global features contained in the frequency domain can also offer a more comprehensive understanding of image patterns. By applying frequency-domain convolution and other attention mechanisms, it enables frequency learning that filters out irrelevant components and focuses on target-specific component range. The feasibility of frequency learning in COD has been demonstrated. Zhong et al. [25] introduced offline frequency cues and proposed a frequency perceptual loss to learn them. Cong et al. [41] used octave convolution to separate high and low frequency features, learning them to obtain coarse localization information. He et al. [42] applied spatial and channel attention to highlight notable frequency parts. Sun et al.[24] further leveraged self-attention for global frequency modeling. Liu et al. [43] proposed to reduce noise in frequency domain by contrastive learning. Zhang et al. [23] employed global average pooling and convolution to generate weights for adaptively learning frequency amplitudes. Previous methods are limited by the locality of convolution, the efficiency of self-attention, and the issue of semantic incoherence when performing global modeling across different frequency components. We propose a frequency-domain sequential scanning strategy to ensure the continuity of components, while using SSMs for efficient adaptive frequency learning. Our method also perceives motion cues through variations in frequency energy, further exploiting frequency information.
@@ III-A Preliminaries
Cross Scan in VMamba. Traditional Mamba focuses on 1D sequence modeling and struggles to handle 2D spatial information in images. To address this, Vision Mamba [31] proposes a bidirectional scanning mechanism that unfolds the image into forward and backward sequences along the horizontal direction. As shown in Fig. 2, Vmamba [32] further introduces 2D selective scan using a cross-scan strategy along four directional paths, corresponding to the vertical, horizontal, and their reverse directions. Each of the four sequences is individually processed through a S6 module for long-range dependency modeling. Subsequently, the sequences are reshaped back into 2D space via cross merge, achieving more comprehensive and effective spatial modeling.
Figure 3: Overview of the proposed Vcamba (left), and RFVSS module (right). $I_{t}$ denotes to the input, which concatenates the windowed sequence containing N frames, $P_{i}$ represents the hierarchical prediction of our model, and $G_{t}$ denotes the ground truth corresponding to each input frame. We adopt VMamba [1] pretrained on ImageNet [2] as encoder, which contains 4 visual state space (VSS) layer. Each VSS Decoder Layer consists of one visual state space block.
@@ III-B Overview
Figure 4: Diagram of the adaptive frequency component enhancement (AFE) module.
@@ III-C Adaptive Frequency Component Enhancement
We introduce frequency features into VCOD to address the limited discriminative capability of spatial features caused by foreground-background similarity. In addition, compared to spatial features, frequency features offer a global perspective for perceiving camouflaged objects. Previous studies [43, 47, 25] have shown that high-frequency components capture fine-grained spatial details while low-frequency components encode structural information. VCOD requires both high-frequency details for subtle edge cues and low-frequency structures for localization. However, frequency components often contain background noise, making it necessary to learn frequency components in an adaptive manner, so that the components corresponding to camouflaged objects can be amplified while those related to background noise are suppressed.
Traditional frequency learning methods employ either convolution or self-attention to dynamically enhance frequency components. However, convolutional approaches are constrained by limited receptive fields and fail to capture global frequency dependencies, while self-attention mechanisms suffer from high computational cost and encounter semantic disparity issues across frequency components during global modeling. Since different frequency components correspond to different semantic levels when transformed back to spatial domain, excessive disparity among them may lead to semantic incoherence, making it hard to interpret their dependencies.
As shown in Fig. 4, we propose a frequency-domain sequential scanning strategy that unfolds the frequency map in sequential order, enabling long-range dependency modeling over the frequency component sequence. Since SSM models sequences by passing historical information through hidden states, the prediction at each time step depends on a continuous memory of past components, especially the most recent one. Thus it makes the learning of current frequency component strongly relies on preceding ones. Our sequential scanning strategy ensures semantic continuity among adjacent components, avoiding inaccurate long-range dependencies when frequency components rely on others with excessive disparity, as it may cause difficulty in mutual understanding.
Specifically, the AFE module transforms spatial features into frequency domain via 2D FFT in a channel-wise manner. In the frequency map, component magnitude correlates with radial distance from center - lowest frequencies at center, highest at corners. Therefore, the concentric circles of different radii represent different frequency bands. Our sequential scanning strategy unfolds frequency components in both low-to-high and high-to-low orders. The low-to-high scanning starts at the center and spirals outward along increasing radii in a clockwise direction with a step size of 1, forming a spiral scanning pattern. The high-to-low sequence is generated by reversing the low-to-high sequence. The bi-directional frequency component sequences are then separately fed into two S6 modules to model long-range dependencies, enabling adaptive enhancement. The enhanced frequency features are then transformed back to spatial domain via iFFT for subsequent motion perception.

### v2

# Mamba-based Spatio-Frequency Motion Perception for Video Camouflaged Object Detection

@@ Abstract
Existing video camouflaged object detection (VCOD) methods primarily rely on spatial appearances for motion perception. However, the high foreground-background similarity in VCOD limits the discriminability of such features (e.g. color and texture). Recent studies demonstrate that frequency features can not only compensate for appearance limitations, but also perceive motion through dynamic variations in spectral energy. Meanwhile, the emerging state space model called Mamba enables efficient motion perception in frame sequences with its linear-time long-sequence modeling capability. Motivated by this, we propose Vcamba, a visual camouflage Mamba based on spatio-frequency motion perception that integrates frequency and spatial features for efficient and accurate VCOD. Specifically, by analyzing the spatial representations of frequency components, we reveal a structural evolution pattern that emerges from the ordered superposition of components. Based on this observation, we propose a unique frequency-domain sequential scanning (FSS) strategy to unfold the spectrum. Utilizing FSS, the adaptive frequency enhancement (AFE) module employs Mamba to model the causal dependencies within sequences, enabling effective frequency learning. Furthermore, we propose a space-based long-range motion perception (SLMP) module and a frequency-based long-range motion perception (FLMP) module to model spatio-temporal and frequency-temporal sequences. Finally, the space and frequency motion fusion module (SFMF) integrates dual-domain features into unified motion representation. Experiments show that Vcamba outperforms state-of-the-art methods across 6 evaluation metrics on 2 datasets with lower computation cost, confirming its superiority. Code is available at: https://github.com/BoydeLi/Vcamba.
@@ I Introduction
However, it is challenging for both humans and machines to accurately localize camouflaged objects based solely on appearances in static scenes. In nature, organisms identify prey or predators by perceiving dynamic changes in the environment, as motion reveals inconsistencies with the static background. This phenomenon aligns with the two-stream hypothesis of neuroscience [24], which indicates that motion is crucial for breaking camouflage. Based on this, existing video camouflaged object detection (VCOD) methods [7, 35, 34, 74] break camouflage by explicitly modeling motion cues in frames like optical flow or implicitly learning them through temporal context.
@@
Although existing methods have made progress, relying solely on spatial appearance features limits performance. The high similarity between foreground and background in texture, color, and pattern results in limited discriminative cues and increased background noise in spatial domain. Recent COD methods [72, 58, 77] have demonstrated the effectiveness of frequency features for enhancing foreground-background distinction. More critically, spatial features are limited by local pixel dependencies and lack global semantic awareness, whereas frequency features inherently provide global representation [8, 54]. Additionally, previous studies [11, 38] have shown that inter-frame motion can be effectively represented by dynamic variations in frequency-domain energy. Therefore, Integrating frequency features into VCOD supplements discriminative cues and enables global motion perception, compensating for spatial feature limitations. As shown in Fig. 1, frequency features enable our model to deliver finer detail segmentation and more accurate global localization.
@@
In addition, most existing COD and VCOD methods [27, 34, 72] are based on convolutional neural networks (CNNs) and Transformers. However, CNN-based methods struggle to capture global features due to local receptive fields, while Transformer-based methods enable global modeling through self-attention, but facing efficiency issues from quadratic complexity. Since VCOD requires both global perception and low computation cost, the emerging state space model (SSM) called Mamba [26] enables global modeling in linear-time, making it an ideal fit for VCOD, which has been extended from natural language processing (NLP) to vision tasks for efficient long-range modeling of images after converting them into 1D representations [80, 44]. It achieves strong performances in downstream visual tasks [44, 23, 53].
However, the application of Vision Mamba in VCOD and its synergy with frequency remain underexplored. Directly employing it for encoding fails to fully realize its potential. The core of VCOD lies in breaking camouflage via motion cues. In essence, motion corresponds to object variations along the temporal sequence, which can be manifested both in spatial features and frequency-domain energy. Thus, inter-frame motion perception fundamentally involves modeling long-range dependencies across frame sequences in both spatial and frequency domains, as the interpretation of the current frame relies on accumulated cues from previous frames. Mamba naturally fits such causal dependency modeling.
Meanwhile, directly leveraging frequency features introduces additional noise. Existing methods [72, 58, 9] have begun exploring frequency modeling in COD, typically using convolution or self-attention to reweight or correlate spectral components, thereby adaptively focusing on target-relevant components. However, these approaches mainly focus on static correlations between components, without explicitly considering the long-range dependencies among them. In fact, by analyzing the spatial representations of spectral components, we discover that they exhibit a structural evolution pattern during ordered superposition, which aligns closely with the progressive process by which human vision and neural networks uncover camouflaged objects. In contrast, SSMs can continuously pass historical information through hidden states, enabling current output to build upon structures formed by preceding components, which corresponds to the superposition process of frequency components. Therefore, through Mamba, we can perceive long-range dependencies in component sequences and uncover latent structural evolution patterns, achieving effective frequency learning.
Based on all these insights, we propose Vcamba, a Mamba-based spatio-frequency motion perception network. Specifically, We employ VMamba [44] as the backbone, given its widespread adoption in downstream tasks [29, 4, 76]. We propose an adaptive frequency enhancement (AFE) module that introduces a novel frequency-domain sequential scanning (FSS) strategy. This strategy unfolds frequency components in both low-to-high and high-to-low orders. By propagating components sequentially, the SSM enables their ordered superposition at the spatial representation level, thereby capturing latent structural evolution patterns and achieving adaptive frequency enhancement. We further introduce a dual-branch framework composed of a space-based long-range motion perception (SLMP) module and a frequency-based long-range motion perception (FLMP) module. The two branches adopt spatio-temporal and frequency-temporal scanning strategies, respectively, organizing frames into spatio-temporal and frequency-temporal sequences. By modeling spatial structural variations and spectral energy changes of targets over time, they extract complementary dual-domain motion cues. Additionally, we analyze the spatial representations of amplitude and phase in camouflaged scenes. As shown in Fig. 1, visualizations indicate that amplitude encodes appearance details, while phase emphasizes structure and contours. According to the Fourier shift theorem, spatial displacement induces linear changes in phase while leaving amplitude unchanged [20]. Therefore, FLMP leverages phase variations to perceive structural motion, effectively suppressing appearance interference. Finally, we propose a spatial and frequency motion fusion (SFMF) module, which aggregates dual-domain motion cues via point-to-point and sequence-to-sequence concatenation for unified motion representation. The main contributions are as follows:
@@
  • We propose a frequency-domain sequential scanning strategy that unfolds FFT-derived frequency components in low-to-high and high-to-low spiral order to achieve adaptive frequency component enhancement.
@@
  • Our Vcamba achieves state-of-the-art performance across six evaluation metrics on two benchmark datasets, and has lower computation cost compared to other VCOD methods, demonstrating the effectiveness and efficiency of our method.
@@ II-A Video Camouflaged Object Detection
Existing COD methods [17, 37, 77, 70] use CNNs and transformers to extract static visual features like color and texture to break camouflage. Fan et al. [17] first introduced COD task and established a two-stage search-and-identify network as a baseline. However, static visual features are easily disrupted by camouflage, making them less discriminative. In contrast, dynamic motion cues across video frames are considered crucial for breaking camouflage. Traditional video methods exploit motion cues [10] or optical flow [32] to perceive camouflaged objects. Cheng et al. [7] argued that the ambiguous and indistinct boundaries of camouflaged objects hinder the accuracy of explicit motion estimation methods such as optical flow. To address this, they proposed a Transformer-based network called SLTNet that implicitly captures motion through long-range modeling. Hui et al. [34] integrated implicit and explicit motion learning into a unified framework, proposing IMEX, which leverages global frame alignment and explicit motion supervision. Then, they further constructed self-prompts to harness the powerful Segment Anything Model [41] for precise segmentation in [35]. Most of the above methods explore motion cues in the spatial domain. However, discriminative features in the spatial domain are limited due to the foreground-background similarity. We further explore frequency feature and integrate it with spatial motion cues to achieve complementary enhancement.
@@ II-B Visual Mamba
Mamba [25] is a general sequence modeling architecture based on the selective State Space Model (SSM). Its SSM effectively captures long-range dependencies, demonstrating linear time complexity relative to sequence length. This addresses the limitations of Transformers in long-range dependency modeling performance and computational efficiency. Mamba was initially proposed in NLP for language sequences modeling [26]. Then, the first visual Mamba backbone, Vision Mamba [80], was introduced. It utilizes a bidirectional scanning method to serialize images into 1D sequences, with the SSM capturing long-range dependencies in the image sequence. Based on this, Liu et al. [44] proposed the more powerful Vmamba, which performs four-directional scanning in both horizontal and vertical directions. Furthermore, to better adapt Visual Mamba to downstream salient object detection, He et al. [29] proposed a saliency Mamba called Samba, which employs a spatial neighboring scanning strategy to preserve the spatial continuity of salient objects. In addition, Mamba has also been widely applied to image classification [23], semantic segmentation [44], and multi-modal learning [53]. However, the application of Mamba in VCOD remains under-explored. Since motion cues are crucial for VCOD and inherently arise from long-range temporal dependencies in frame sequences, we employ Mamba to model both spatio-temporal and frequency-temporal sequences. Furthermore, we also introduce a novel frequency-domain scanning strategy that enables Mamba to perform adaptive frequency learning.
@@ II-C Frequency Learning
Frequency learning is widely used in signal analysis to reveal the time-frequency characteristics and energy distribution patterns of signals. The rich global features contained in the frequency domain can also offer a more comprehensive understanding of image patterns. By applying frequency-domain convolution and other attention mechanisms, it enables frequency learning that filters out irrelevant components and focuses on target-specific component range. The feasibility of frequency learning in COD has been demonstrated. Zhong et al. [77] introduced offline frequency cues and proposed a frequency perceptual loss to learn them. Cong et al. [9] used octave convolution to separate high and low frequency features, learning them to obtain coarse localization information. He et al. [28] applied spatial and channel attention to highlight notable frequency parts. Sun et al.[58] further leveraged self-attention for global frequency modeling. Liu et al. [45] proposed to reduce noise in frequency domain by contrastive learning. Zhang et al. [72] employed global average pooling and convolution to generate weights for adaptively learning frequency amplitudes. Previous methods rely on convolution or self-attention for frequency learning, while overlooking the long-range dependencies within frequency component sequences. By analyzing the spatial representations of frequency components, we uncover a latent structural evolution pattern emerging from their ordered superposition, which aligns with the process of object discovery in COD. It can be effectively captured by Mamba’s causal modeling to facilitate more powerful frequency learning.
@@ III-A Preliminaries
Cross Scan in VMamba. Traditional Mamba focuses on 1D sequence modeling and struggles to handle 2D spatial information in images. To address this, Vision Mamba [80] proposes a bidirectional scanning mechanism that unfolds the image into forward and backward sequences along the horizontal direction. As shown in Fig. 2, VMamba [44] introduces 2D selective scan using a cross-scan strategy along four directional paths, corresponding to the vertical, horizontal and their reverse directions. Each of the four sequences is processed through a S6 module for long-range dependency modeling. Subsequently, the sequences are reshaped back into 2D space via cross merge, achieving more comprehensive and effective spatial modeling.
Phase-based Motion Perception. According to the Fourier shift theorem [48], a spatial translation introduces a linear phase shift in the frequency domain while leaving the amplitude unchanged, and phase is more sensitive and robust than amplitude to contrast variations and affine deformations [20]. Based on this principle, Hommos et al. [31] showed that phase variations directly correspond to object motion, making phase a viable representation for motion modeling. Phase information has also been widely exploited in motion representation and video processing [61, 78, 52].
Specifically, for an image $f(x,y)$, its 2D Fourier transform is denoted as $F(u,v)$:
$F(u,v)=\iint f(x,y)\,e^{-j2\pi(ux+vy)}\,dx\,dy,$ (4)
after spatial translation by $(\Delta x,\Delta y)$,
$g(x,y)=f(x-\Delta x,\,y-\Delta y),$ (5)
the transformed spectrum becomes $G(u,v)$:
$G(u,v)=F(u,v)\,e^{-j2\pi(u\Delta x+v\Delta y)},$ (6)
specifically, the amplitude is obtained by applying the modulus operation $|\cdot|$ to $G(u,v)$, while the phase is computed by taking the argument (i.e., the angle) of the complex spectrum using the arctangent function $\angle(\cdot)$:
$|G(u,v)|=\left|F(u,v)\,e^{-j2\pi(u\Delta x+v\Delta y)}\right|=|F(u,v)|,$ (7)
$\angle G(u,v)=\angle F(u,v)-2\pi(u\Delta x+v\Delta y),$
Where $|G(u,v)|$ denotes the amplitude of $G(u,v)$, the translation introduces a complex exponential term whose modulus equals one, leaving the amplitude spectrum unchanged. $\angle G(u,v)$ denotes the phase of $G(u,v)$, which undergoes a linear shift with respect to the phase of $F(u,v)$.
Figure 3: Overview of the proposed Vcamba. $I_{t}$ denotes to the input, which concatenates the windowed sequence containing N frames, $P_{i}$ represents the hierarchical prediction of our model, and $G_{t}$ denotes the ground truth corresponding to each input frame. Each VSS Decoder Layer consists of one VSSBlock.
@@ III-B Overview
Figure 4: Diagram of the adaptive frequency component enhancement (AFE) module and the frequency-domain sequential scanning (FSS) strategy.
@@ III-C Adaptive Frequency Component Enhancement
Due to the inherent concealment and ambiguity, spatial modeling of camouflaged objects based on local pixel dependencies often lacks sufficient discriminability and global context. Frequency-domain representations provide a complementary global view, where different spectral components encode spatial representations at distinct granularities. Previous studies [45, 66, 77] have shown that high-frequency components capture fine-grained spatial details while low-frequency components encode structural information. Since camouflaged object cues and background noise coexist across components, effective COD requires both low-frequency structural modeling and high-frequency detail perception. Existing frequency-based COD methods [72, 58, 9] have explored adaptive modulation of components or inter-component correlations. However, these methods mainly focus on component-wise reweighting or correlation modeling, without perceiving long-range dependencies or exploiting the underlying dependency pattern across components.
Since long-range dependencies are inherently order-sensitive, and frequency learning aims to modulate the spatial representations of components, we seek to mine the latent dependency order of spatial representations. Inspired by U-Net [55], where object representations are progressively formed via hierarchical feature superposition, the ordered superposition of components naturally aligns with this structural evolution process. Consequently, we propose the frequency-domain sequential scanning (FSS) strategy.
FSS unfolds the spectrum into continuous low-to-high and high-to-low sequences. As shown in Fig. 4, spatial features are transformed into the frequency domain via 2D FFT. The low-to-high sequence starts from the spectral center and spirals outward along increasing radii in a clockwise manner, while the high-to-low sequence is obtained by reversing this order. Since SSMs model causal sequences by passing historical information through hidden states, each output depends on preceding inputs in a causal manner. This mechanism closely mirrors the ordered superposition process of frequency components. The two bi-directional sequences are thus processed by parallel S6 modules to perceive long-range dependencies and exploit structural evolutionary patterns, resulting in enhanced frequency representations. The enhanced frequency features are then transformed back to the spatial domain via iFFT for subsequent motion perception.
Figure 5: Diagram of the structural evolution pattern emerging from the ordered superposition of frequency components. The low-to-high and high-to-low sequences are constructed by superposing components unfolded through the FSS. We provide a schematic illustration of the superposition process for each sequence, along with their corresponding spatial features.
In order to illustrate the rationale of FSS, we visualize the ordered superposition of frequency components in Fig. 5. By sequentially superposing frequency components and reconstructing their spatial representations, one can see that the low-to-high sequence simulates the natural visual process from coarse perception to local refinement, which closely matches the coarse-to-fine paradigm of U-Net. Conversely, one can see that the high-to-low sequence demonstrates an evolution from fine details to holistic structures. Both sequences embody a structural evolution pattern with causal relations, and by modeling such long-range dependencies, SSMs enable more effective frequency learning, thereby facilitating accurate object detection.